OpenStudy (anonymous):

Suppose f and g are continuous functions such that g(9) = 5 and lim x → 9 [3f(x) + f(x)g(x)] = 24. Find f(9).

OpenStudy (anonymous):

plug in g(9) to your equation: [3f(9)+f(9)*g(9)]=24 3f(9)+f(9)*5=24. Now f(9) is just like any normal variable so you can add the two f(9)'s together

OpenStudy (anonymous):

so f(9)[3+5]=24. f(9)=24/8=3

OpenStudy (anonymous):

thanks!